# Facilitating Automated Online Consensus Building through Parallel Thinking

### v1

@@ Abstract.
Consensus building is inherently challenging due to the diverse opinions held by stakeholders. Effective facilitation is crucial to support the consensus building process and enable efficient group decision making. However, the effectiveness of facilitation is often constrained by human factors such as limited experience and scalability. In this research, we propose a Parallel Thinking-based Facilitation Agent (PTFA) that facilitates online, text-based consensus building processes. The PTFA automatically collects textual posts and leverages large language models (LLMs) to perform all of the six distinct roles of the well-established Six Thinking Hats technique in parallel thinking. To illustrate the potential of PTFA, a pilot study was carried out and PTFA’s ability in idea generation, emotional probing, and deeper analysis of ideas was demonstrated. Furthermore, a comprehensive dataset that contains not only the conversational content among the participants but also between the participants and the agent is constructed for future study.
collective intelligence, consensus building, automated facilitation agent, parallel thinking, six hats
$c$s: Human-centered computing Empirical studies in collaborative and social computing
1. Introduction
@@
Collective intelligence, the emergent capability of groups to solve problems and make decisions collaboratively, has become a cornerstone of effective decision making in diverse domains, from business strategy to public policy (Becker et al., 2017). By integrating the knowledge, creativity, and perspectives of multiple individuals, collective intelligence can produce outcomes that surpass those achievable by individuals working in isolation. However, harnessing this potential often requires skilled facilitation to guide discussions, manage conflicts, and ensure equitable participation. To address the need for effective group consensus building facilitation, artificial intelligence (AI)-based approaches leveraging machine learning and intelligent agents have been proposed to support tasks such as content generation, balancing participation, and accumulating facilitation knowledge (Ito et al., 2022; Hadfi et al., 2023; Gu et al., 2021). In particular, the advent of generative AI and large language models (LLMs) has demonstrated significant potential in enhancing facilitation tasks, including collaborative brainstorming (Nomura et al., 2024), generating and refining statements (Tessler et al., 2024) and mitigating conflicts (Ma et al., 2024a). Despite these advances, achieving systematic automated facilitation support remains challenging. It requires abilities such as sophisticated contextual understanding, maintaining the natural flow of discussions, and striking a delicate balance between structured guidance and flexibility to avoid rigidity. Moreover, the significant energy cost of training LLMs and the popularity of well-known LLMs like ChatGPT makes it desirable to use pre-trained LLMs as well as necessary to investigate the capabilities of existing LLMs for such a specific application.
@@
To bridge the gap between specific facilitation task support and systematic automated facilitation, we propose a novel Parallel Thinking-based Facilitation Agent (PTFA) designed to support online group consensus building. This conversational agent actively participates in text-based online discussions, guiding groups through structured, context-aware interactions to foster mutual understanding and agreement. It integrates the principles of parallel thinking, a structured methodology for systematically exploring multiple perspectives, with the advanced capabilities of large language models (LLMs) (Göçmen and Coşkun, 2019; Wan et al., 2024). Specifically, PTFA is based on the Six Thinking Hats technique, where each hat represents a role of an effective facilitator, e.g., the blue hat represents the role of managing the discussion and the green hat represents fostering creativity. The roles and suitable discussion interactions are selected automatically and dynamically based on the contributions of the participants. By dynamically adapting to the flow of discussions, PTFA adjusts its facilitation role in response to the evolving context, generating targeted facilitation inputs to enrich the discussion.
However, new and up-coming applications in this domain, like PTFA, will likely rely on readily-available LLMs, which require fine-tuned instructions based on their characteristics. The availability and popularity of certain LLMs, like ChatGPT and Gemini, means that, at least in the short term, they are likely to be used for LLM-based applications, including tools that are intended to support discussion facilitation. Hence, we based our PTFA implementation on a ChatGPT system prompt customization that uses the ChatGPT API to generate facilitator responses. By utilizing this approach we are able to provide insights that this approach has the potential to improve AI-based automatic facilitation, but needs specific tuning to the timeliness, phasing of discussion and consistency to provide an effective facilitator and a practical tool.
To evaluate PTFA and the general-purpose LLM approach, we developed an online text-based discussion platform, serving as an interactive interface for participants to gather and conduct discussions. We conducted a pilot study involving 16 discussion groups and 48 participants in an online environment collecting a rich dataset on participant interaction with the PTFA which has allowed us to establish guidelines on pursuing AI-based consensus-building platform development in idea generation, emotional probing and deeper argument analysis as well as challenges such as automated phase management and dynamic response timing control. Furthermore, this dataset will be public available when the full paper gets accepted.
2. Related Work
2.1. Group Decision Support with Automated Facilitation
@@
Group decision making benefits significantly from facilitation, which can occur during, before, or after group discussions (Bostrom et al., 1993; Beranek et al., 1993; Zaraté et al., 2013). Traditionally, human facilitators have been essential in guiding discussions, requiring expertise in both the subject matter and the language of communication (Gimpel et al., 2024). However, the shortage of skilled facilitators creates a gap that modern automated facilitation systems can help address (Gu et al., 2021). Most existing automated facilitation approaches rely on rule-based and static reasoning systems (Gu et al., 2021). While these systems provide structure, their lack of adaptability limits their effectiveness in dynamic group interactions. Recently, LLMs have gained prominence as tools for brainstorming and consensus building, offering a dynamic and adaptive alternative. LLMs engage in active collaboration with human users, overcoming the rigidity of traditional systems (Noack et al., 2025; Wan et al., 2024; Joosten et al., 2024; Bouschery et al., 2023). LLMs have demonstrated promise in several aspects of group decision making. They facilitate agreement building within groups (Tessler et al., 2024; Shin et al., 2022), enhance the quality of discourse (Argyle et al., 2023), and contribute to better decision-making outcomes (Ma et al., 2024a). Additionally, LLMs can foster diversity of opinions (Kim et al., 2020) and improve inclusivity, such as by increasing women’s participation in discussions (Hadfi et al., 2023).
Despite these advances, the assessment of LLMs as generalized facilitators, capable of performing multiple tasks simultaneously while actively participating in conversations, remains underexplored. This paper aims to evaluate the effectiveness of LLMs in such a role, specifically through the lens of the Six Thinking Hats methodology (De Bono, 2017). Furthermore, we assess the quality of deliberations from the user’s perspective, a dimension that has been identified as lacking sufficient research (Shortall et al., 2022).
2.2. Facilitation with the Six Thinking Hats Method
@@
The Six Thinking Hats framework is a well-established method for enhancing group decision-making and problem-solving. It involves six metaphorical hats, each representing a distinct mode of thinking. By focusing on one type of thinking at a time, the method reduces confusion and improves efficiency in group discussions (De Bono, 2017). The Six Thinking Hats methodology has been employed in experimental settings with encouraging results, particularly for creative brainstorming (Göçmen and Coşkun, 2019; Fender and Stickney, 2017). Previous studies have explored the use of this framework in facilitating human reflection on LLM-generated outputs, where participants evaluate and critique ideas rather than having LLMs actively assume the roles of the six hats (Zha et al., 2024).
@@
In contrast, our approach assigns each LLM agent the role of one of the Six Thinking Hats to facilitate real-time discussions with human participants. While prior work has explored similar roles for LLMs in agent-to-agent discussions (Lu et al., 2024; Mushtaq et al., 2025), to the best of our knowledge this study is the first to evaluate their effectiveness in facilitating human group interactions through this method. The Six Thinking Hats approach can lead to higher-quality argumentation and foster critical thinking (Ekahitanond, 2018). However, it may also increase the cognitive load for participants, requiring careful consideration of its implementation (Ma et al., 2024b). By leveraging LLMs to assume these roles, we aim to mitigate the cognitive burden on human participants while enhancing the overall quality and depth of discussions. Our study is the first to assess the impact of using LLMs to actively embody the Six Thinking Hats roles in real-time group discussions, providing new insights into their potential for automated facilitation and collaborative decision-making.
3. Methodology
Overall: The discussion platform is developed using the Discourse forum system  [^1] https://www.discourse.org/, integrated with LLMs to facilitate structured conversations based on the Six Thinking Hats methodology (De Bono, 2017). The platform assists users in reaching consensus through structured interactions. The system consists of several key components, including a Discourse-based web interface where participants engage in discussions, a back-end database for storing discussion logs and participant interactions, and the OpenAI  [^2] https://openai.com/Assistants API, which provides real-time responses from the six thinking hats agents. WebSockets are used to enable real-time communication between the front-end and back-end components, while third-party plugins enhance forum functionalities such as logging and analytics.
In terms of technical implementation, the system utilizes the Discourse API to retrieve users’ post content and associated metadata (including timestamps, user IDs, and topic IDs), and the OpenAI Assistants API to generate responses based on the Six Thinking Hats framework. Discussion logs are securely stored in a PostgreSQL database, and participants are assigned unique identifiers to maintain anonymity. Security measures include HTTPS encryption to secure communications and role-based access control to manage permissions.
[…]
system
Figure 1. The Framework of PTFA.
@@ 3. Methodology
Parallel Thinking-based Facilitation Agent Implementation: The facilitation agents in the system are implemented using OpenAI’s ChatGPT assistant API, which enables real-time responses to guide structured discussions based on the Six Thinking Hats framework (shown in Fig.1). Each thinking hat is represented by a distinct LLM agent that provides contextually appropriate responses. The White Hat agent provides factual data without interpretation, the Red Hat agent encourages the sharing of feelings and gut reactions, and the Black Hat agent identifies potential risks and downsides. The Yellow Hat agent highlights positive aspects and potential benefits, the Green Hat agent encourages innovative ideas and alternatives, and the Blue Hat agent directs the conversation and ensures structure.
@@
The implementation of LLM-based facilitation agents requires careful prompt engineering at both a macro and micro level. At a macro level, prompts should be designed to provide comprehensive context by outlining the discussion background, defining the agent’s role, and specifying the overall objectives. This includes instructing the agent to guide participants towards consensus, foster balanced discussions, and adhere to structured communication principles. At the micro level, prompts should incorporate specific instructions that allow the agent to dynamically assess the current state of the conversation and intervene appropriately. These prompts should include mechanisms to determine when intervention is necessary, such as identifying signs of topic drift, unresolved conflicts, or opportunities for further elaboration. The prompts also define response styles, ensuring that the agent’s interventions are concise, relevant, and aligned with the discussion phase.
@@
Prompt engineering for these LLM-based facilitation agents is designed to ensure that each agent’s responses align with its designated role while maintaining clarity and conciseness. The primary prompts provide general guidance based on the respective thinking mode, whereas situational prompts encourage specific interventions tailored to the evolving discussion context. For instance, the White Hat prompts focus on factual accuracy and verification, while the Green Hat encourages brainstorming new ideas. The agents are triggered based on predefined timing intervals. Timing intervals are set based on the progression of the discussion phase, with interventions occurring at 30-second intervals to ensure continuous engagement without overwhelming the participants. If the system detects a lack of diversity in input or prolonged inactivity, the agents are prompted to re-engage users with targeted suggestions or requests for further elaboration.
4. User Study
4.1. Participants & Study Setting
@@
To evaluate the efficacy of the proposed PTFA in supporting online consensus building, we held text-based real-world online discussions. Before embarking on this study, we secured ethical approval from the Faculty Ethics Committee at XX University (hidden for anonymization), guaranteeing adherence to established guidelines for ethical conduct in research involving human subjects. A total of 48 participants were recruited through lecture announcements with faculty permission, email invitations via department mailing lists, and posters placed in university common areas. 16 discussion groups were formed and each discussion group consisted of 3 participants, who were randomly selected while ensuring a balance in gender and English proficiency levels in each group. Detailed participant characteristics are shown in Table 1 in the appendix. Specifically, each group included at least two different genders and participants with at least two different levels of English proficiency. The experiments were conducted in university tutorial rooms equipped with laptops that connected to the online discussion platform we developed. Figure 3 and Figure 4 in the appendix show examples of experiment scenarios and the screenshot of the online discussion platform, respectively. Each discussion group participated in two online discussions, each featuring a different combination of discussion topics and facilitation models, detailed setting is shown in Table 1 in the appendix, and followed by a short facilitator evaluation survey.
@@
Each discussion session lasted for 20 minutes. Two topics related to daily life were selected as the discussion subjects. Topic0 was ”Please decide one activity that you would like to do together” and Topic1 was ”Please decide one film that you would like to watch together”. Additionally, two facilitation models were designed for comparison. Facilitation Model 0 represented a traditional facilitation model, in which three facilitation messages were generated based on the discussion’s progress. Specific facilitation messages can be found in the appendix. In contrast, Facilitation Model 1 was developed using the proposed PTFA approach, introducing an alternative facilitation method for comparison.
4.2. Results

### v2

@@ Abstract
Consensus building is inherently challenging due to the diverse opinions held by stakeholders. Effective facilitation is crucial to support the consensus building process and enable efficient group decision making. However, the effectiveness of facilitation is often constrained by human factors such as limited experience and scalability. In this research, we propose a Parallel Thinking-based Facilitation Agent (PTFA) that facilitates online, text-based consensus building processes. The PTFA automatically collects real-time textual input and leverages large language models (LLMs) to perform all six distinct roles of the well-established Six Thinking Hats technique in parallel thinking. To illustrate the potential of the agent, a pilot study was conducted, demonstrating its capabilities in idea generation, emotional probing, and deeper analysis of idea quality. Additionally, future open research challenges such as optimizing scheduling and managing behaviors in divergent phase are identified. Furthermore, a comprehensive dataset that contains not only the conversational content among the participants but also between the participants and the agent is constructed for future study.
1 Introduction
@@
Collective intelligence, the emergent capability of groups to solve problems and make decisions collaboratively, has become a cornerstone of effective decision making in diverse domains, from business strategy to public policy [2]. By integrating the knowledge, creativity, and perspectives of multiple individuals, collective intelligence can produce outcomes that surpass those achievable by individuals working in isolation. However, harnessing this potential often requires skilled facilitation aimed at supporting decision making, such as guiding the process, promoting constructive interaction and inclusivity, and managing content in a structured manner [9]. To address the need for effective group consensus building facilitation, artificial intelligence (AI)-based approaches leveraging machine learning and intelligent agents have been proposed to support tasks such as content generation, balancing participation, and accumulating facilitation knowledge [13, 12, 10]. In particular, the advent of generative AI and large language models (LLMs) has demonstrated significant potential in enhancing specific facilitation tasks, including collaborative brainstorming [22], generating and refining statements [25] and mitigating conflicts [17]. Despite recent advances, achieving automated systematic facilitation support remains challenging, particularly because it involves managing multiple facilitation tasks concurrently and managing the relationships between those tasks in dynamic decision-making processes. For example, systematic facilitation requires the ability to understand the context of the decision-making process, determine how to promote participation, and adapt accordingly based on the identified characteristics and needs of the participants. Additionally, it must determine the appropriate timing for executing these facilitation tasks. This requires abilities such as sophisticated contextual understanding, maintaining the natural flow of discussions, and striking a delicate balance between structured guidance and flexibility to avoid rigidity.
@@
To bridge the gap between specific facilitation task support and systematic automated facilitation, we propose a novel Parallel Thinking-based Facilitation Agent (PTFA) designed to support online group consensus building. This conversational agent actively participates in the group decision-making process, guiding stakeholders through structured, context-aware interactions to foster mutual understanding and consensus. To address multiple facilitation tasks concurrently, PTFA integrates the principles of parallel thinking, a structured methodology for systematically exploring multiple perspectives, with the advanced capabilities of LLMs [11, 26]. Specifically, PTFA is based on the Six Thinking Hats technique, which is a successful method for finding consensus [6]. Each hat represents a role of an effective facilitator, e.g., the blue hat represents the role of managing the discussion and the green hat represents fostering creativity. The roles and suitable discussion interactions are selected automatically and dynamically based on the contributions of the participants. To evaluate PTFA, we developed an online text-based discussion platform, serving as an interactive interface for participants to gather and conduct discussions. We conducted a pilot study involving 16 discussion groups and 48 participants in an online environment collecting a rich dataset on participant interaction with the PTFA which has allowed us to establish guidelines on pursuing AI-based consensus-building platform development in idea generation, emotional probing and deeper argument analysis as well as challenges such as automated phase management and dynamic response timing control. To summarize, the main contributions of this paper are as follows:
  • We propose a novel agent-based framework that implements a multi-role facilitation model powered by an LLM, enabling automated and systematic support for group decision-making.
  • We implement the proposed approach in an online, free-text decision-making scenario and conduct real-world pilot experiments to evaluate its potential in idea generation, emotional probing, and deeper analysis of idea quality and find limitations in optimizing scheduling and managing behaviors as a fully automated facilitator.
  • We construct a novel dataset rich in conversational interactions, capturing dialogues both among participants and between participants and the agent.
The remainder of this paper is organized as follows. Section 2 reviews related work on automated facilitation, conversational agents, and parallel thinking methodologies. Section 3 introduces the overall framework of the proposed agent-based facilitation system. Section 4 details the implementation and experimental setup, and presents the results of the pilot experiments, along with a discussion of their implications, limitations, and potential directions for future work. Finally, Section 5 concludes the paper.
2 Related Work
2.1 Group Decision Support with Automated Facilitation
@@
Group decision making benefits significantly from facilitation, which can occur during, before, or after group discussions [4, 3, 27]. Traditionally, human facilitators have been essential in guiding discussions, requiring expertise in both the subject matter and the language of communication [9]. However, the shortage of skilled facilitators creates a gap that modern automated facilitation systems are increasingly positioned to address [10]. Most existing automated facilitation approaches rely on rule-based and static reasoning systems [10]. While these systems provide structure, their lack of adaptability limits their effectiveness in dynamic group interactions. Recently, LLMs have gained prominence as tools for brainstorming and consensus building, offering a dynamic and adaptive alternative. In contrast to rigid systems, LLMs engage in active collaboration with human users, supporting more natural facilitation [21, 26, 14, 5]. LLMs have demonstrated promise in several aspects of group decision-making. They facilitate agreement building within groups [25, 23], enhance the quality of discourse [1], and contribute to better decision-making outcomes [17]. Additionally, LLMs can foster diversity of opinions [15] and improve inclusivity, such as by increasing women’s participation in discussions [12].
However, prior LLM-based facilitation approaches tend to focus on technical optimization, or surface level coordination [19], overlooking research in other fields that offer structured methods for enhancing group discussion. We argue that effective automated facilitation should be grounded in established methodologies that are designed to support reflective, balanced, and goal-oriented discourse. To this end, this paper explores the role of LLMS as generalized facilitators, able to perform different tasks and roles, through the lens of the Six Thinking Hats methodology [6]. This structured approach to parallel thinking is well-established in the organizational literature but has not been utilized in automated facilitation. Furthermore, we assess the quality of deliberations from the user’s perspective, a dimension that has been identified as lacking sufficient research [24].
2.2 Facilitation with the Six Thinking Hats Method
@@
The Six Thinking Hats framework is a well-established method for enhancing group decision-making and problem-solving. It involves six metaphorical hats, each representing a distinct mode of thinking. By focusing on one type of thinking at a time, the method reduces confusion and improves efficiency in group discussions [6]. The Six Thinking Hats methodology has been employed in experimental settings with encouraging results, particularly for creative brainstorming [11, 8]. Previous studies have explored the use of this framework in facilitating human reflection on LLM-generated outputs, where participants evaluate and critique ideas rather than having LLMs actively assume the roles of the six hats [28].
@@
In contrast, our approach assigns each LLM agent the role of one of the Six Thinking Hats to facilitate real-time discussions with human participants. While prior work has explored similar roles for LLMs in agent-to-agent discussions [16, 20], to the best of our knowledge, this study is the first to evaluate their effectiveness in facilitating human group interactions through this method. The Six Thinking Hats approach can lead to higher-quality argumentation and foster critical thinking [7]. However, it may also increase the cognitive load for participants, requiring careful consideration of its implementation [18]. By leveraging LLMs to assume these roles, we aim to mitigate the cognitive burden on human participants while enhancing the overall quality and depth of discussions. Our study is the first to assess the impact of using LLMs to actively embody the Six Thinking Hats roles in real-time group discussions, providing new insights into their potential for automated facilitation and collaborative decision-making.
3 Methodology
Overall Implementation. To support the PTFA framework, we developed an online discussion platform based on the Discourse forum system [^1] https://www.discourse.org/, enhanced with large language model (LLM) integration. The platform enables structured, real-time text-based discussions, with LLM agents acting as facilitators according to the Six Thinking Hats methodology. The system architecture consists of a web-based interface for user participation, a backend database to store user-generated content and metadata, and an integration with the OpenAI Assistants API [^2] https://openai.com/ for dynamic response generation. Communication between frontend and backend components is handled using WebSockets to ensure real-time responsiveness.
To manage discussions securely and effectively, each participant is assigned a unique anonymous identifier, and all discussion data is stored in a PostgreSQL database. The platform employs HTTPS encryption for secure communication and role-based access control to restrict system access. In addition, third-party plugins are used to support logging, moderation, and analytics, providing additional functionality for monitoring and evaluating discussion quality. The LLM agents are triggered based on predefined time intervals or inactivity detection, ensuring responsive but non-intrusive facilitation throughout the conversation.
Figure 1: The framework of PTFA.
@@ 3 Methodology
Parallel Thinking-based Facilitation Agent Implementation. The facilitation agents in the system are implemented using OpenAI’s ChatGPT assistant API, which enables real-time responses to guide structured discussions based on the Six Thinking Hats framework (shown in Fig.1). Each thinking hat is represented by a distinct LLM agent that provides contextually appropriate responses.
@@
The implementation of LLM-based facilitation agents requires careful prompt engineering at both a macro and micro level. At a macro level, prompts should be designed to provide comprehensive context by outlining the discussion background, defining the agent’s role, and specifying the overall objectives. This includes instructing the agent to guide participants towards consensus, foster balanced discussions, and adhere to structured communication principles. At the micro level, prompts should incorporate specific instructions that allow the agent to dynamically assess the current state of the conversation and intervene appropriately. These prompts should include mechanisms to determine when intervention is necessary, such as identifying signs of topic drift, unresolved conflicts, or opportunities for further elaboration. The prompts also define response styles, ensuring that the agent’s interventions are concise, relevant, and aligned with the discussion phase. The prompt engineering for each of the six hat agents is detailed below:
  • White Hat (Facts and Information): This agent embodies objective thinking. Its prompt instructs it to focus solely on data and facts without interpretation, grounding the conversation in reality. It is designed to intervene when the discussion requires objective information or when participants make unsupported statements. An example of a situational prompt is: "Could you clarify the exact figures or facts related to this issue? Here’s what we know so far: [insert relevant data]."
  • Red Hat (Emotions and Intuition): Representing feelings and gut reactions, this agent allows for the expression of emotion without justification. The primary prompt directs it to provide a short, intuitive perspective when the conversation touches on personal feelings. A situational prompt for this agent is: "This feels like an emotional moment. How are we feeling about this issue right now?"
  • Black Hat (Critical Thinking and Risk): This agent provides a critical, cautious perspective. Its core instruction is to identify potential risks, weaknesses, and flaws that the group might be overlooking, ensuring negative outcomes are considered. An example of its situational prompt is: "Have we considered the potential downsides? Here’s a risk we might be overlooking: [insert risk]."
  • Yellow Hat (Optimism and Benefits): The Yellow Hat agent focuses on optimism and positive thinking. It is prompted to offer a constructive view by highlighting the benefits and potential advantages of ideas, especially when the discussion needs a boost of optimism. A corresponding situational prompt is: "Looking at the bright side, this idea offers some exciting opportunities we shouldn’t overlook."
  • Green Hat (Creativity and Alternatives): This agent encourages creative and innovative thinking. It is prompted to intervene with new ideas or alternative approaches when the conversation becomes stalled or could benefit from novel solutions. An example of a creative situational prompt is: "What if we approached this from a different angle? Here’s an idea to consider: [insert new idea]."
  • Blue Hat (Process and Control): The Blue Hat agent manages the thinking process itself. Its instructions are to provide structure and direction, organize the conversation, and ensure the discussion stays on track and that all perspectives are considered. A situational prompt for managing the process is: "It seems like we’re getting off track. Maybe we should focus on this key point: [insert key point]."
@@
Prompt engineering for these LLM-based facilitation agents is designed to ensure that each agent’s responses align with its designated role while maintaining clarity and conciseness. The primary prompts provide general guidance based on the respective thinking mode, whereas situational prompts encourage specific interventions tailored to the evolving discussion context. For instance, the White Hat prompts focus on factual accuracy and verification, while the Green Hat encourages brainstorming new ideas. The agents are triggered based on predefined timing intervals. Timing intervals are set based on the progression of the discussion phase, with interventions to ensure continuous engagement without overwhelming the participants. If the system detects a lack of diversity in input or prolonged inactivity, the agents are prompted to re-engage users with targeted suggestions or requests for further elaboration.
4 User Study
4.1 Participants & Study Setting
@@
To evaluate the efficacy of the proposed PTFA in supporting online consensus building, we held text-based real-world online discussions. Before embarking on this study, we secured ethical approval from the Faculty Ethics Committee at University of Southampton, guaranteeing adherence to established guidelines for ethical conduct in research involving human subjects. A total of 48 participants were recruited through lecture announcements with faculty permission, email invitations via department mailing lists, and posters placed in university common areas. 16 discussion groups were formed and each discussion group consisted of 3 participants, who were randomly selected while ensuring a balance in gender and English proficiency levels in each group. Specifically, each group included at least two different genders and participants with at least two different levels of English proficiency. The experiments were conducted in university tutorial rooms equipped with laptops that connected to the online discussion platform we developed. Figure 2 shows the screenshot of the online discussion platform. Each discussion group participated in two online discussions, each featuring a different combination of discussion topics and facilitation models and followed by a short facilitator evaluation survey. The purpose of the survey was to evaluate the participants’ perception of the facilitator’s effectiveness, overall user experience and the degree to which consensus was achieved. By comparing the responses between the facilitator models, we aim to evaluate the practical impact of PTFA on group discussion dynamics and outcomes. Through this process we identify areas of improvement for LLM-assisted facilitation.
Figure 2: Screenshot of online discussion platform
@@
Each discussion session lasted for 20 minutes. To alleviate barriers to participation, two topics related to daily life were selected as discussion subjects. Topic0 was "Please decide one activity that you would like to do together" and Topic1 was "Please decide one film that you would like to watch together". Additionally, two facilitation models were designed for comparison. Facilitation Model 0 represented a traditional facilitation model in which three facilitation messages were generated, each after a certain amount of elapsed time:
  • “Hi all, our goal today is to reach a consensus on the question posed at the end of the discussion. Please start by generating ideas.” was posted at the beginning of the discussion,
  • “You have already discussed it for 10 mins. This is a good time for you to reconsider the ideas that you have already had.” was posted 10 minutes after the discussion began, and
  • “There are only 3 minutes left, if you haven’t reached a consensus yet, please make a decision as soon as possible.” was posted 17 minutes after the discussion started.
In contrast, Facilitation Model 1 was developed using the proposed PTFA approach, introducing an alternative facilitation method for comparison. Intervention decisions are made at 30-second intervals, taking group size into account.
4.2 Results
Our study consisted of 48 participants, with 3 in each of the 16 groups. Each group participated in two discussion giving a dataset of 32 discussions containing 16,656 words in 1,669 posts, of which the facilitator contributed 3,459 words in 217 posts. Each post has a corresponding timestamp and, for the LLM-generated posts, which hat was used to generate the text. Participants were almost entirely aged between 18 and 34, with a near even split between males and females. Most participants (68.8%) considered themselves a fluent or native speaker of english, with only 2.1% considering themselves a beginner, the least proficient category. A full breakdown of the participant demographics is presented in Figure 3.
(a) The share of the participants by their ages.
(b) The share of the participants by their genders.
(c) The share of the participants by their proficiencies in English.
Figure 3: The characteristics of the participants in terms of age, gender and English proficiency.